# On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

### v1

@@ 5 Numerical Verification
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$A_{i}$ $=$ $randn(m,n)\in{\mathbb{R}}^{m\times n},\;\textrm{for}\;i=0:3,$
$B_{i}$ $=$ $randn(m,d)\in{\mathbb{R}}^{m\times d},\;\textrm{for}\;i=0:3,$
$C_{i}$ $=$ $randn(p,n)\in{\mathbb{R}}^{p\times n},\;\textrm{for}\;i=0:3,$
$D_{i}$ $=$ $randn(p,d)\in{\mathbb{R}}^{p\times d},\;\textrm{for}\;i=0:3.$
Take $m=30t$, $n=10t$, $p=2t$, and $d=2$. Here, $t$ is an arbitrary number. We apply Algorithm 2 to determine the real solution of the RBLSE problem. Let $X_{RL}$ be the real solution of the RBLSE problem
$\min_{X,R_{r}}\left\lVert R_{r}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{r},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{RL}-(B+R_{r})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{RL}-D\right\rVert_{F}$. In Table 3, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $2.9063\times 10^{-15}$ $3.0851\times 10^{-15}$
$3$ $2.5053\times 10^{-15}$ $5.5184\times 10^{-15}$
$5$ $3.7915\times 10^{-15}$ $1.0949\times 10^{-14}$
$7$ $3.9204\times 10^{-15}$ $1.3185\times 10^{-14}$
$9$ $5.2026\times 10^{-15}$ $1.7247\times 10^{-14}$
@@
Table 3: Computational accuracy of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 3 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-14}$. This indicates that Algorithm 2 is highly effective in determining the real solution for the RBLSE problem (29).
Example 5.4.
In Example 5.3, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the real solution $X_{RL}$ of the RBLSE problem (29). These perturbations are measured normwise by the smallest $\epsilon$, as in (38). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$. Table 4 presents the computed exact relative forward errors and the corresponding upper bound $U_{RL}$ (calculated using Equation 40) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$
$1$ $7.3638\times 10^{-12}$ $2.3814\times 10^{-11}$ $8.0339\times 10^{-10}$
$7.3677\times 10^{-10}$ $2.1476\times 10^{-9}$ $8.0381\times 10^{-8}$
$7.9614\times 10^{-8}$ $7.6845\times 10^{-7}$ $8.6858\times 10^{-6}$
$3$ $7.1742\times 10^{-14}$ $2.6216\times 10^{-13}$ $9.0524\times 10^{-12}$
$8.1702\times 10^{-11}$ $1.5855\times 10^{-10}$ $1.0309\times 10^{-8}$
$7.1758\times 10^{-9}$ $2.6438\times 10^{-8}$ $9.0544\times 10^{-7}$
$5$ $7.2567\times 10^{-11}$ $2.4717\times 10^{-10}$ $1.0430\times 10^{-8}$
$7.2735\times 10^{-9}$ $2.5130\times 10^{-8}$ $1.0454\times 10^{-6}$
$8.7538\times 10^{-7}$ $1.8374\times 10^{-6}$ $1.2581\times 10^{-4}$
Table 4: Comparison of relative forward errors and their upper bounds for the real solution of a perturbed RBLSE problem (29)
Table 4 shows that the exact relative forward errors of the real solution to the RBLSE problem (29) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{RL}$ for the relative forward error.
Example 5.5.
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 10}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{3\times 10}$, and $X_{0}=rand(10,3)+rand(10,3)\textit{{i}}\in{\mathbb{C}}^{10\times 3}$. Let
$M_{i}$ $=$ $rand(m,10)+rand(m,10)\textit{{i}}\in{\mathbb{C}}^{m\times 10},\;\textrm{for}\;i=1,2$
$R_{i}$ $=$ $rand(3,10)+rand(3,10)\textit{{i}}\in{\mathbb{C}}^{3\times 10},\;\textrm{for}\;i=1,2.$
Take $B=AX_{0}$ and $D=CX_{0}$. Thus, $X_{0}$ is the complex solution to the RBLSE problem (11). To measure the performance of the proposed technique for the RBLSE problem (11), Algorithm 1 is employed to determine the complex solution $X$. Let the error be $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 5.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$100$ $1.3154\times 10^{-14}$
$200$ $6.2150\times 10^{-15}$
$300$ $6.5603\times 10^{-15}$
$400$ $6.1485\times 10^{-15}$
$500$ $3.2441\times 10^{-15}$
Table 5: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 5 shows that the error $\epsilon$ between the complex solution derived from Algorithm 1 and the corresponding true solution to the RBLSE problem (11) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 1 in computing the complex solution to the RBLSE problem (11).
Example 5.6.
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 50}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{10\times 50}$, and $X_{0}=randn(50,30)\in{\mathbb{R}}^{50\times 30}$. Let
$A_{i}$ $=$ $randn(m,50)\in{\mathbb{R}}^{m\times 50},\;\textrm{for}\;i=0:3$
$C_{i}$ $=$ $randn(10,50)\in{\mathbb{R}}^{10\times 50},\;\textrm{for}\;i=0:3.$
Take $B=AX_{0}$ and $D=CX_{0}$. Clearly, $X_{0}$ is the real solution to the RBLSE problem (29). To measure the performance of the proposed technique for the RBLSE problem (29), Algorithm 2 is employed to determine the real solution $X$. Define the error as $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 6.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$1000$ $3.8948\times 10^{-14}$
$2000$ $4.4732\times 10^{-14}$
$3000$ $4.1257\times 10^{-14}$
$4000$ $3.7531\times 10^{-14}$
$5000$ $4.6532\times 10^{-14}$
Table 6: Computational accuracies of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 6 shows that the error $\epsilon$ between the real solution derived from Algorithm 2 and the corresponding true solution to the RBLSE problem (29) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 2 in computing the real solution to the RBLSE problem (29).
@@ 6 Conclusions
We have presented an algebraic method for solving the RBLSE problem by transforming it into complex and real LSE problems. This transformation leverages the complex and real representations of reduced biquaternion matrices, enabling efficient computation of both complex and real solutions to the RBLSE problem. Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the robustness and accuracy of our method in addressing the RBLSE problem.
@@ 7 Declarations
The first author played a crucial role in overseeing the project and providing valuable guidance. Neha Bhadala made significant contributions at various stages of the research, including conceptualization, methodology design, software implementation, validation, and formal analysis. Additionally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.

### v2

@@ 9 Compute $X$ $O(8n^{2}d)$
We define the problem parameters as follows:
Let $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $C\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$ as defined in Example 5.1. For $i=0,1$, let $X_{i}=rand(n,d)\in{\mathbb{R}}^{n\times d}$. Using these components, we define:
$X_{R}=X_{0}\in{\mathbb{R}}^{n\times d},\quad X_{C}=X_{0}+X_{1}\textit{{i}}\in{\mathbb{C}}^{n\times d}$
Here, the parameters $m$, $p$, $n$, and $d$ remain consistent with Example 5.1 and depend on the variable $t$.
For the real solution, we set $B=AX_{R}$ and $D=CX_{R}$, which ensures that $X_{R}$ is the exact real solution of the RBLSE problem. Algorithm 1 is employed to compute the approximate real solution $\widehat{X}_{R}$. The accuracy of the computed solution is evaluated using the Frobenius norm error:
$\epsilon_{R}=\|X_{R}-\widehat{X}_{R}\|_{F}.$
For the complex solution, we set $B=AX_{C}$ and $D=CX_{C}$, making $X_{C}$ the exact complex solution of the RBLSE problem. Algorithm 2 is used to compute the approximate complex solution $\widehat{X}_{C}$, and the error is defined as:
$\epsilon_{C}=\|X_{C}-\widehat{X}_{C}\|_{F}.$
The variable $t$ determines the dimensions of the matrices $A$ and $C$, and the errors $\epsilon_{R}$ and $\epsilon_{C}$ are computed for various values of $t$. The results are summarized in Table 5, which shows the relationship between $t$ and the computational accuracy of the algorithms.
$t$ $\epsilon_{R}=\left\lVert X_{R}-\widehat{X}_{R}\right\rVert_{F}$ $\epsilon_{C}=\left\lVert X_{C}-\widehat{X}_{C}\right\rVert_{F}$
$1$ $7.0655\times 10^{-15}$ $3.5401\times 10^{-15}$
$3$ $1.3490\times 10^{-14}$ $1.0711\times 10^{-15}$
$5$ $2.3275\times 10^{-14}$ $2.2491\times 10^{-14}$
$7$ $5.4268\times 10^{-14}$ $3.7398\times 10^{-14}$
$9$ $5.8569\times 10^{-14}$ $5.2257\times 10^{-14}$
@@
Table 5: Computational accuracy of Algorithms 1 and 2 for computing the real and complex solutions of the RBLSE problem.
Table 5 demonstrates that the errors $\epsilon_{R}$ and $\epsilon_{C}$ are consistently below $10^{-14}$ for all tested values of $t$. These results confirm the high accuracy of the proposed algorithms for computing the real and complex solutions of the RBLSE problem, showcasing their effectiveness and precision across varying problem sizes.
@@ 6 Conclusions
We have presented an algebraic method for solving the RBLSE problem by transforming it into real and complex LSE problems. This transformation leverages the real and complex representations of reduced biquaternion matrices, enabling efficient computation of real and complex solutions to the RBLSE problem. Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the accuracy and reliability of our method in addressing the RBLSE problem. Future research will focus on exploring their applications in image and signal processing.
@@ 7 Declarations
The first author played a crucial role in overseeing the project and providing valuable guidance. Neha Bhadala made significant contributions at various stages of the research, including conceptu-alization, methodology design, software implementation, validation, and formal analysis. Additio-nally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.
Conflict of interest
This work does not have any conflicts of interest.